The wave function of a particle is πœ“(π‘₯) = { π΄π‘π‘œπ‘ ( 2πœ‹π‘₯ 𝐿 ) π‘“π‘œπ‘Ÿ βˆ’ 𝐿 4 ≀ π‘₯ ≀ 𝐿 4 0 elsewhere i) Determine the normalization constant A. ii) What is the probability that the particle will be found between x= 0 and x = L/6 if we measured its position? iii) Find the expectation values for the operators x, p, and p 2 .


1
Expert's answer
2021-06-16T14:25:41-0400
"\int_{-L/4}^{L/4}\psi^2(x)dx = 1"

Substituting the wavefunction, obtain:



"\int_{-L/4}^{L/4}A^2\cos^2(2\pi x/L)dx = 1\\"

Taking the integral, find:



"A^2\left( \dfrac{L}{2\pi}\right)\dfrac{\pi }{2} = 1\\
A = \dfrac{2}{\sqrt{L}}"

The probability of finding the particle between 0 and L/8 is:



"\int_{0}^{L/8}\psi^2(x)dx =\int_{0}^{L/8}A^2\cos^2(2\pi x/L)dx = \dfrac{1}{4} + \dfrac{1}{2\pi} \approx 0.409"

Answer.Β "A = \dfrac{2}{\sqrt{L}},p = 0.409".